OpenStudy (anonymous):

b1= r * (Sy/Sx) How do I get the number for r (correlation) given the averages of X and Y and their standard deviations?

OpenStudy (anonymous):

@asnaseer @bahrom7893 @iHelp @jim_thompson5910 @karatechopper

OpenStudy (anonymous):

Statistics? fun. \[r = sxy/((sx)(sy))\] where sx and sy are you standard devaitions and sxy is:

OpenStudy (anonymous):

Which is the average of your means.

OpenStudy (anonymous):

im still a little confused. can you tell me how they got b1 based on the data presented?

OpenStudy (anonymous):

@.Sam. @eyust707 @satellite73

OpenStudy (anonymous):

missing a lot of the page

OpenStudy (anonymous):

\[S_{xy}=\frac{1}{n-1}\sum_{i=1}^n(x_1-\overline{x})(y_1-\overline{y})\] i can understand, we just need these numbers

OpenStudy (anonymous):

well then ill upload it as an excel file rather than a PDF because PDF gets choppy

OpenStudy (anonymous):